OpenStudy (pink.1d):

For an angle Θ with the point (12, -5) on its terminating side, what is the value of cosine?

OpenStudy (michele_laino):

by definition of polar coordinates, we can write: \[\left\{ \begin{gathered} 12 = r\cos \theta \hfill \\ - 5 = r\sin \theta \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

squaring both formulas above, and summing them each other, we get: \[144 + 25 = {r^2}\] So what is r?

OpenStudy (pink.1d):

r is 13

OpenStudy (michele_laino):

that's right!

OpenStudy (triciaal):